Q: A body weighing 2 kg tied at the end of string 1 meter long and rotating in a vertical circular path at a speed of 6 meter per second. What is the tension at the string at top and at bottom?
1
Expert's answer
2012-12-28T04:38:12-0500
T = mg + ma
T - the tension at the string
a - the acceleration of the body
a = v^2/r
On the top:
T = ma - mg = m(v^2/r - g) = 2*(6^2/1-10) = 52 H
On the bottom:
T = ma + mg = m(a + g) = 2*(36+10) = 92 H